# Containerization in Multi-Cloud Environment: Roles, Strategies, Challenges, and Solutions for Effective Implementation

### v1

@@ 2.2.2. Snowballing
In Phase 2, we utilized the snowballing technique, as described in (Wohlin, 2014), to examine references within primary studies for identifying additional relevant studies. This strategy was augmented by forward snowballing, where we gathered studies that cited the selected studies, and backward snowballing, which involved using references within the selected studies. Notably, we encountered several dozen studies during the forward and backward snowballing that had been excluded in primary search. This phase resulted in the addition of 8 more studies, bringing the final count to 86.
@@ 2.3. Data Extraction and Analysis
Data Extraction: The data extraction form was designed based on predefined data items (see Table 5) that were formulated to address the RQs specified in Table 1. To ensure the reliability of the extracted data items, the pilot data extraction was conducted on ten studies by the first author, and all the other authors assessed the extracted data. Subsequently, the first author employed a revised set of data items (e.g., D11, D12, D16), determined after evaluating the extracted data items, for the formal data extraction from the selected studies. To mitigate the personal bias and ambiguity, all authors engaged in discussions regarding the extracted data. The data items labeled as D1 to D11 present a summary of the demographics of the primary studies selected, while D12 to D26 are specifically employed to address RQ1 to RQ6. A concise description of each data item is also presented in Table 5. Finally, Google Sheets were used to record and further analysis of the extracted data.
Data Analysis: We employed a descriptive statistics to analyze the quantitative data from data items D1, D5, D7, D8, and D9. For the remainder of the data, which primarily consist of qualitative free-text descriptions (e.g., study aim, roles of containers, challenges, and solutions), we conducted a thematic analysis in accordance with the guidelines outlined in (Braun and Clarke, 2006). Our thematic analysis process consists of the following steps:
@@
D14 Automation Challenges Challenges in automation discussed in the study RQ4
D15 Automation Solutions Solutions for automation challenges discussed in the study RQ5
@@
D20 Security Challenges Challenges in security discussed in the study RQ4
D21 Security Solutions Solutions for security challenges discussed in the study RQ5
@@
The process of obtaining research themes executed in this SMS is shown in Figure 2. Two researchers participated in the process in order to reduce personal bias. The most important activity of this process is brainstorming sessions that were mainly conducted while reviewing, defining, and naming the research themes. In these sessions, both researchers discussed and validated the research themes found.
Concerning data item D9 (testing approaches) and D10 (testing challenges), we used the open coding and constant comparison techniques from Grounded Theory (Glaser and Strauss, 1967) to analyze the qualitative data extracted from the selected studies.
Finally, we provided a replication package (Waseem et al., 2024) containing the results of each phase of the study selection process (e.g., Phase 1, Phase 2) and detailed results (e.g., Contributions, Patterns, QAs, Challenges, Solutions) for verification and validation purposes of this SMS.
@@ 3. Results
This section presents the results of the SMS,derived from analyzing data from chosen studies. Section 3.1 provide a summary of the demographic data, classifications, and the mapping of research themes identified in the studies. Section 3.2 presents the results concerning container roles and implementation strategies. Section 3.3 reports on the patterns and strategies employed in containerized-based applications in multi-cloud environment. Section 3.4 presents the QAs and related tactics for containerized applications in multi-cloud environment. Section 3.5 presents the security challenges and solution framework, while Section 3.6 presents the automation challenges and solution framework. Section 3.7 presents the deployment challenges and solution framework, and Section 3.8 presents the monitoring challenges and solution framework. Lastly, the tools and frameworks are reported in Section 3.9.
@@ 3.1. Demographics, Research Aims, and Contributions
Studies Yearly Distribution: Figure 3-a illustrates the annual distribution of studies found in eight databases from January 2013 to March 2023. After applying our selection criteria, the final dataset includes studies published from 2014 to 2022. We found no primary studies on containerization in multi-cloud environments meeting our criteria in the first three months of 2023. The bars displays yearly study counts, with a trend showing the overall direction. A peak in 2017, followed by a decrease and occasional fluctuations, suggests an early spike in interest then more targeted research.
@@
Publishers Distribution: The pie chart presented in Figure 3-b illustrates the distribution of identified studies among different publishers. The largest portion of the studies, accounting for 72.09%, is published by IEEE, indicating its significant contribution to the field. Scopus follows with 11.62%, while Engineering Village, ACM, and Springer Link represent smaller portions at 6.97%, 5.81%, and 3.48%, respectively. The chart emphasizes the dominance of IEEE in the identified studies, suggesting its importance as a primary source of relevant literature.
@@
Study Types: Figure 3-a provides a breakdown of the types of the selected studies. It is apparent that conferences dominate as the most common venue, comprising a substantial 72.09% of the total publications. Journals, which are often considered as venues for more mature and rigorously peer-reviewed research, make up 19.76%. This indicates a strong preference for disseminating information through conferences, which may be due to the faster publication process and opportunities for networking and collaboration. Symposiums and workshops account for 6.97% and 1.16% respectively, suggesting that these are less common avenues, possibly due to their more specialized or focused nature.
@@
Authors’ Affiliations: Figure 3-c illustrates the distribution of authors’ affiliations in the collected publications. A staggering 93.02% of the affiliations are associated with academia. This overwhelmingly high percentage emphasizes the central role that academic institutions play in the production and dissemination of research in the field. Industry affiliations represent a mere 4.65%, which could imply that industry professionals may be more focused on practical applications and might not publish frequently in the venues considered. Mixed affiliations, at 2.32%, indicate a small proportion of collaborations between academia and industry. This low rate of cross-affiliation collaboration may suggest that there is potential for increased synergy between academic and industrial entities in future research endeavors.
@@
Mapping of Research Themes: The thematic analysis of selected studies aims on containerization in multi-cloud environment, as illustrated in Figure 4, indicating various significant insights.The prominent research themes identified include “Scalability and High Availability” with 20 studies, and three other themes: ’Security and Privacy”m “Multi-Cloud Container Monitoring and Adaptation”, and “Performance and Optimization”, each attracting attention in 15 studies. This highlights that the core areas of interest in multi-cloud environment are proficient container management, scalability, and performance optimization. “Security and Privacy’ also stands as a considerable theme, as demonstrated by the 15 studies focusing on areas such as secure container handling, data security, and blockchain applications. This emphasizes the necessity of protecting data and applications in the context of containers spreading across multi-cloud environment.
Moreover, it is noted that several studies encapsulate more than one theme. For instance, study S19 is catalogued under both ‘Multi-cloud Orchestration and Management’ and ‘Application Deployment in Multi-cloud Container’, signifying that it encompasses aspects of both Kubernetes and microservices architecture.
In essence, the exploration into containerization in multi-cloud environment is substantially inclined towards, scalability, performance optimization security, and orchestration, with several studies showcasing inter-thematic relevance. This is suggestive of the diverse challenges and opportunities present in deploying containers within multi-cloud contexts.
Major Contributions of the Selected Studies: Table 6 provides an overview of the contributions made by the selected studies. The contributions are primarily categorized into five main categories and 22 subcategories (see the Contributions sheet in (Waseem et al., 2024)), through thematic analysis. The selected studies primarily focused on proposing and validating reference architectures, security approaches, cloud management and interoperability frameworks, orchestration and deployment approaches, as well as approaches and tools for optimization and deployment of containerized applications in multi-cloud environment.
The studies related to the reference architecture category mainly contributed to the understanding and design of architectures specifically tailored for multi-cloud containerization. They aimed to address the complexities associated with containerization in multi-cloud environment. On the other hand, the studies in the security approaches category discuss concerns related to cloud analysis, security and performance, cybersecurity, and access control within multi-cloud contexts.
Figure 4. Selected studies research aims in containerization for multi-cloud environment.
The studies on cloud management and interoperability frameworks primarily aim to improve resource and service management, monitoring, and deployment in multi-cloud environment. Their goal is to enhance the performance and efficiency of managing resources and services across multiple clouds. Meanwhile, the research in the area of orchestration and deployment approaches introduces methods for bench-marking, optimizing, and simulating container orchestration in multi-cloud settings. The objective of these studies is to streamline the process of deploying and orchestrating containers within multi-cloud environment.
Lastly, the research on optimization and deployment approaches and tools focuses on improving container deployment and creating efficient management solutions for multi-cloud environment. These studies underscore the importance of optimizing resource allocation and developing practical management tools for use in multi-cloud settings. Overall, these studies contribute to proposing reference architectures, security approaches, cloud management, and interoperability frameworks. These studies also offer strategies and tools for orchestration, deployment operations, and optimization in multi-cloud containerization.
@@
Reference Architectures and Models for Cloud General Cloud Architectures S26, S33, S35, S42, S76
Fault Tolerance & Self-Healing S12, S55, S67, S74
Migration & Testing Architecture S17, S79
Domain-Specific Language S80, S82
Performance Modeling S28, S62
Optimization & Reliability Model S50, S66
IoT & Service Mesh Architecture S05, S59
Security Approaches Cloud Analysis Approach S18, S24, S37, S65, S69, S86
Security and Performance Solution S10, S23, S38, S75, S77
Cybersecurity & Access Control S11, S30, S34, S71, S73
Security Approach S06, S25, S40, S83
Cloud Management and Interoperability Framework Resource & Service Management Framework S48, S49, S47, S15, S58
Monitoring & Deployment Framework S19, S45, S46, S08
Interoperability Framework S03, S81
Evaluation & Modeling Framework S20, S85
Orchestration and Deployment Approach Benchmarking & Optimization Approach S01, S02, S41, S51, S68
Service Management Approach S07, S14, S29
Simulation & Integration Approach S27, S31, S56
Storage & Services Solution S64, S75
Optimization and Deployment Approaches and Tool Optimization & Scheduling Approach S22, S32, S36, S39
Multi-Cloud Platform S09, S16, S43, S60, S72
Deployment Approach and Tool S21, S52, S57, S61

### v2

@@ 2.3.2. Snowballing
In Phase 2, we utilized the snowballing technique, as described in (Wohlin, 2014), to examine references within 110 primary studies to identify additional relevant studies. This strategy was augmented by forward snowballing, where we gathered studies that cited the selected studies, and backward snowballing, which involved using references within the selected studies. Notably, we encountered several dozen studies during the forward and backward snowballing that had been excluded in primary search. This phase resulted in the addition of 11 more studies, bringing the final count to 121.
2.3.3. Quality Assessment of Selected Studies
We assessed the quality of the selected studies based on a set of predefined criteria, including relevance to multi-cloud containerization, clarity of research design, empirical validation (e.g., experiments, case studies), reporting quality (e.g., discussion of limitations and challenges), publication in peer-reviewed venues, and the generalizability of findings to multi-cloud environment. Each study was evaluated using a scoring system where criteria were rated as 0 (not satisfied), 0.5 (partially satisfied), or 1 (fully satisfied), with a maximum total score of 6. Studies scoring below 3 were flagged for further review and discussion but retained to ensure alignment with the study’s objectives. This quality assessment was conducted after the study selection process to ensure consistency and thoroughness across all 121 studies. Detailed results are available in the replication package (Waseem et al., 2024), specifically in the sheet titled “QA of Selected Studies”.
@@ 2.4. Data Extraction and Analysis
Data Extraction: The data extraction form was designed based on predefined data items (see Table 5) that were formulated to address the RQs specified in Table 1. To ensure the reliability of the extracted data items, the pilot data extraction was conducted on ten studies by the first author, and all the other authors assessed the extracted data. Subsequently, the first author employed a revised set of data items (e.g., D11, D12, D16), determined after evaluating the extracted data items, for the formal data extraction from the selected studies. To mitigate the personal bias and ambiguity, all authors engaged in discussions regarding the extracted data. The data items labeled as D1 to D8 present a summary of the demographics of the primary studies selected, while D9 to D22 are specifically employed to address RQ1 to RQ6. A concise description of each data item is presented in Table 5. Finally, Google Sheets were used to record and further analyze the extracted data.
Data Analysis: We employed descriptive statistics to analyze the quantitative data from data items D1, D5, D7, D8, and D9. For the remainder of the data, which primarily consist of qualitative free-text descriptions (e.g., study aim, roles of containers, challenges, and solutions), we conducted a thematic analysis in accordance with the guidelines outlined in (Braun and Clarke, 2006). Our thematic analysis process consists of the following steps:
@@
D14 Automation Challenges Challenges in automation discussed in the study RQ4.2
D15 Automation Solutions Solutions for automation challenges discussed in the study RQ5.2
@@
D20 Security Challenges Challenges in security discussed in the study RQ4,1
D21 Security Solutions Solutions for security challenges discussed in the study RQ5.1
@@
Two researchers participated in the process to reduce personal bias. The most important activity of this process is brainstorming sessions that were mainly conducted while reviewing, defining, and naming the research themes. In these sessions, both researchers discussed and validated the research themes found.
Concerning data item D9 (Roles of Containers) and D10 (Container Implementation strategy), we used the open coding and constant comparison techniques from Grounded Theory (Glaser and Strauss, 1967) to analyze the qualitative data extracted from the selected studies.
Finally, we provided a replication package (Waseem et al., 2024) with the results of each phase of the study selection process (e.g., Phase 1, Phase 2) and detailed results (e.g., Contributions, Patterns, QAs, Challenges, Solutions) for verification and validation purposes of this SMS.
@@ 3. Results
This section presents the results of the SMS, derived from analyzing data from chosen studies. Section 3.1 provides a summary of the demographic data, classifications, and the mapping of research themes identified in the studies. Section 3.2 presents the results concerning container roles and implementation strategies. Section 3.3 reports on the patterns and strategies employed in container-based applications in multi-cloud environment. Section 3.4 presents the QAs and related tactics for containerized applications in multi-cloud environment. Section 3.5 presents the security challenges and solutions framework, while Section 3.6 presents the automation challenges and solutions framework. Section 3.7 presents the deployment challenges and solutions framework, and Section 3.8 presents the monitoring challenges and solutions framework. Finally, the tools and frameworks are reported in Section 15.
@@ 3.1. Demographics, Research Aims, and Contributions
Yearly Distribution of Studies: Figure 3-a illustrates the annual distribution of studies retrieved from eight databases between January 2013 and July 2024. We found only two primary studies on containerization in multi-cloud environment that meet our criteria in 2013. The bars display yearly study counts and trends. A peak in 2017, followed by a decrease and occasional fluctuations, suggests an early spike in interest followed by more targeted research. Notably, the reduction in publication numbers in 2020 and 2021 may be partially attributed to the global disruptions caused by the COVID-19 pandemic, which affected research outputs across various fields. However, there is a notable increase in studies in 2023 and continuing into 2024. This could be due to delayed research projects getting back on track and more funding in tech sectors after the pandemic, showing a renewed focus on improving containerization in multi-cloud environment
@@
Publishers Distribution: Figure 3-b shows the distribution of studies by publisher. IEEE accounts for 65.29%, highlighting its major contribution. Scopus follows with 11.57%, while Engineering Village, ACM, and Springer Link represent smaller portions at 12.40%, 7.44%, and 2.48%, respectively. The chart emphasizes the dominance of IEEE in the identified studies, suggesting its importance as a primary source of relevant literature. This distribution also suggests to researchers where to primarily search and publish on containerization in multi-cloud environment.
@@
Study Types: Figure 3-a provides a breakdown of the types of selected studies. It is apparent that conferences dominate as the most common venue, comprising 62.5% of the total publications. Journals, often considered venues for mature and rigorously peer-reviewed research, make up 32.5%. This indicates a preference for conferences, likely due to faster publication and networking opportunities Symposiums and workshops represent for 5.0% and 0.83% respectively, suggesting that these are less common avenues, possibly due to their more specialized or focused nature.
@@
Authors’ Affiliations: Figure 3-c illustrates the distribution of authors’ affiliations in collected publications. 90.90% of the affiliations are associated with academia. This highlights the central role of academic institutions in producing and disseminating research. implying that industry professionals focus more on practical applications. The low industry affiliation may reflect the SMS’s focus on academic databases, where industry participation is limited, indicating typical academic database trends and suggesting broader industry inclusion in future studies. Mixed affiliations, at 4.95%, indicate a small proportion of collaborations between academia and industry. This low rate of cross-affiliation collaboration may suggest that there is potential for increased synergy between academic and industrial entities in future research endeavors.
Classification of Research Theme: The identified research themes and their notable trends are illustrated in Figure 4 and detailed below. Figure 4-(a) presents the taxonomical classification of existing research. This taxonomy systematically identifies, names, and represents various topics based on their similarities or distinctions. Figure 4-(b) depicts the trends and temporal distribution of the themes over the years (2013–2024) based on the studies identified in our review.
@@
  • Taxonomy of Research Themes: The prominent research themes identified include “Scalability and High Availability” with 20 studies, and three other themes: “Security and Privacy” in “Multi-Cloud Container Monitoring and Adaptation”, and “Performance and Optimization”, each attracting attention in 15 studies. This highlights that the core areas of interest in multi-cloud environment are proficient container management, scalability, and performance optimization. “Security and Privacy” also stands as a considerable theme, as demonstrated by the 15 studies focusing on areas such as secure container handling, data security, and blockchain applications. This emphasizes the necessity of protecting data and applications in the context of containers spreading across multi-cloud environment. Moreover, it is noted that several studies encapsulate more than one theme. For instance, study S19 is catalogued under both ‘Multi-cloud Orchestration and Management’ and ‘Application Deployment in Multi-cloud Container’, signifying that it encompasses aspects of both Kubernetes and microservices architecture.
  • Trends and Temporal Distribution of Themes: To highlight prominent trends and their progression over the years (2013–2024), we adopted the thematic and sub-thematic representation from Figure 4(a) and illustrated them via a timeline and relative distribution of each theme in Figure 4-(b). Specifically, Figure 4-(b) demonstrates how certain research themes have progressed and matured (i.e., evolved as topics) over the years. For instance, the theme multi-cloud orchestration and management is addressed in a total of 18 studies, accounting for approximately 15% of the total reviewed studies. Our analysis indicates that one of the earliest research efforts on multi-cloud orchestration, published during 2014–2015, focused on the component-based and model-driven adaptation of applications for multi-cloud containers [S1, S20, S8]. Following this initial phase of application adaptation for multi-cloud orchestration, subsequent research explored techniques for benchmarking and enabling resource-efficient application orchestration [S2, S41, S39]. More recent studies, published between 2022 and 2024, have investigated a variety of topics, including but not limited to DevOps-based development [S27] and monitoring orchestrated applications in multi-cloud environments [S39, S87]. Our review revealed no published studies on the theme of multi-cloud orchestration and management during 2013, 2017, or 2021. Two possible reasons for this are: (i) our literature search did not identify any relevant studies (see Section 2.3.1 - primary search), or (ii) the studies from those years did not meet the qualitative evaluation criteria (see Section 2.3.3 - quality assessment) for inclusion in the review.
Figure 4. Selected Studies: a) Thematic Classification and b) Trends Over the Years
Major Contributions of the Selected Studies: Table 6 provides an overview of the contributions of the selected studies. The contributions are primarily categorized into five main categories and 22 subcategories (see the Contributions sheet in (Waseem et al., 2024)), through thematic analysis. The selected studies mainly focus on proposing and validating security solutions, reference architectures, cloud management frameworks, interoperability, orchestration, and tools for optimizing and deploying containerized applications in multi-cloud environment. Notably, the largest contributions come from the Security Approaches category, representing 27.25% of the total studies, which highlights the significant attention to cybersecurity & access control and cloud analysis approaches. Within this category, the Cybersecurity and Access Control subcategory alone accounts for 9.09%, emphasizing the importance of securing multi-cloud environment. The studies related to the Reference Architectures and Models for Cloud category, comprising 21.42% of the studies, focus on understanding and designing architectures specifically tailored for multi-cloud containerization. Within this category, General Cloud Architectures (4.92%) and Fault Tolerance and Self-Healing (4.13%) are key subcategories. These subcategories address critical challenges such as ensuring resilience and performance in cloud-native architectures. The studies on Optimization and Deployment Approaches and Tools category accounts for 18.16% of the studies, indicating a strong research focus on optimizing container deployment and resource scheduling. Within this category, Optimization & Scheduling Approaches (8.26%) represent the second-largest subcategory across all studies, highlighting the practical need for effective resource allocation in multi-cloud systems. Finally, Cloud Management and Interoperability Frameworks, accounting for 12.38%, primarily aim to improve resource utilization, service management, monitoring, and deployment in multi-cloud environments. This includes contributions from Resource and Service Management Frameworks, which represent 4.95% of the studies.
@@
Reference Architectures and Models for Cloud General Cloud Architectures (6, 4.92%) (Watanabe et al., 2022; Sami and Mourad, 2020; Carpio et al., 2020; Baresi et al., 2016; Pop et al., 2016; Alonso et al., 2023a)
IoT & Service Mesh Architecture (5, 4.13%) (Gattobigio et al., 2022; Rodigari et al., 2021; Merlino et al., 2024; Hossain et al., 2023; Solayman and Qasha, 2023)
Fault Tolerance & Self-Healing (5, 4.13%) (Bisht and Kaur, 2022; Bolivar et al., 2018; Tang, 2021; Mfula and Nurminen, 2018; Nagarajan et al., 2023)
Migration & Testing Architecture (3, 2.47%) (Slominski et al., 2015; Serhiienko and Spillner, 2018; Roy and Kalafatis, 2023)
Optimization & Reliability Model (3, 2.47%) (Bhamare et al., 2015; Aral et al., 2020; Huang and Wang, 2023)
Domain-Specific Language (2, 1.65%) (Quint and Kratzke, 2018; Ferry et al., 2013)
Performance Modeling (2, 1.65%) (Barna et al., 2017; Rahman and Lama, 2019)
Security Approaches Cybersecurity & Access Control (11, 9.09%) (He et al., 2022; Sheridan et al., 2017; Rios et al., 2017; Graupner et al., 2015a; Guechi and Maamri, [n. d.])
Cloud Analysis Approaches (7, 5.78%) (Pai and Kumar, 2021; Pahl, 2015; Dimitrijevic et al., 2019; Zeck and Bouroudjian, 2017; Kiss, 2018; Gundu et al., 2020)
Security and Performance Solution (6, 4.95%) (Sabbioni et al., 2020; Wang et al., 2020; Pandey et al., 2022; Ranjbar et al., 2017)
Security Approaches (5, 4.13%) (Deng et al., 2022; Gao et al., 2017; Bélair et al., 2019; Wahab et al., 2016)
Provisioning & Restoration Approach (4, 3.30%) (Pellegrini et al., 2017; Lazuka et al., 2022; Wang et al., 2019; Kritikos et al., 2019)
Cloud Management and Interoperability Framework Resource & Service Management Framework (6, 4.95%) (Saif et al., 2023; Harwalkar et al., 2019b; Libardi et al., 2014; Pham and Pham, 2015; Sitaram et al., 2018; Barletta et al., 2024)
Monitoring & Deployment Framework (5, 4.13%) (Alhamazani et al., 2014; Trihinas et al., 2015; Fadda et al., 2019; Pham et al., 2015; Horchulhack et al., 2024; Marques et al., 2024)
Evaluation & Modeling Framework (4, 3.30%) (Ferry et al., 2014; Bou Ghantous and Gill, 2021; Tundo et al., 2024; Ramesh et al., 2023)
Interoperability Framework (2, 1.65%) (Hua et al., 2020; Challita et al., 2017)
Orchestration and Deployment Approach Benchmarking & Optimization Approach (6, 4.13%) (Almeida et al., 2014; Jha et al., 2019; Shi et al., 2020; Buyya and Barreto, 2015; Guerrero et al., 2018a; Petrovski and Gusev, 2024)
Service Management Approach (4, 3.30%) (Zhang et al., 2021; Chen et al., 2022; Gupta et al., 2021; Benmerar et al., 2023)
Simulation & Integration Approach (3, 2.47%) (Alonso et al., 2019; Mittal and Risco-Martín, 2017; del Castillo et al., 2013)
Storage & Services Solution (3, 2.47%) (Harwalkar et al., 2019a; Shen et al., 2017; Orzechowski et al., 2023)
Optimization and Deployment Approaches and Tool Optimization & Scheduling Approach (10, 8.26%) (Wong et al., 2019; Al-shammari and Alwan, 2018; Lorenzo et al., 2019; Altran et al., 2022; da Silva et al., 2023; Dogani et al., 2023; Pandi et al., 2023; Sofia et al., 2023; Nguyen and Chien, 2023)
Multi-Cloud Platform (6, 4.95%) (Verdugo et al., 2017; Ouyang et al., 2022; Dreibholz et al., 2019; Janarthanan et al., 2018; Alaluna et al., 2017; Wong et al., 2019; Al-shammari and Alwan, 2018; Lorenzo et al., 2019; Altran et al., 2022; Santos et al., 2023)
Deployment Approach and Tool (6, 4.95%) (Cui et al., 2016; Casola et al., 2017; Tricomi et al., 2017; Di Modica et al., 2019; Liu and Qin, 2024; Tiwari and Sharma, 2023)